OpenStudy (anonymous):

length of a rectangle is 4ft longer than its width. if the perimeter of the rectangle is 40 ft, find its area.

OpenStudy (anonymous):

$$96 \text{ feet}$$

OpenStudy (anonymous):

If the width of a rectangle is 'x' feet and then the length of the rectangle is 4 ft longer, then the length is 'x+4' feet If the perimeter is 40ft we know that (2*x)+2*(4+x)=40 so 2x+8+2x=40 x=8ft So width = 8ft length=12ft Area = length * width = 8*12=96 ft^2